# Towards Imitation Learning in Real World Unstructured Social Mini-Games in Pedestrian Crowds

### v1

@@ IV-A Inverse Reinforcement Learning [43, 40]
The objective in MaxEnt IRL is to adjust the weights of the cost function such that the expected feature counts under the derived policy match the feature counts of the expert, while also maximizing the entropy of the policy. The cost function is iteratively adjusted, typically using gradient-based methods, to ensure that the expected feature counts from the policy induced by the current reward function align with those from the expert demonstrations. Note that since the MaxEnt IRL algorithm traditionally applies to a single agent, Problem (1) is modified to finding $\theta^{i,*}$ that minimizes
$\left\lVert\mathbb{E}_{\Gamma^{i}_{\theta^{i}}\sim\mathcal{D}_{\theta^{i}}}\Phi^{i}\left(\Gamma^{i}_{\theta^{i}}\right)-\mathbb{E}_{\widehat{\Gamma}^{i}\sim\mathcal{\widehat{D}}}\Phi^{i}\left(\widehat{\Gamma}^{i}\right)\right\rVert$ (2)
There are several reasons why the MaxEnt IRL has been selected as a representative baseline in this study. First, the maximum entropy principle ensures a unique solution to the reward recovery problem, mitigating ambiguities inherent in traditional IRL. Further, resultant policies are inherently stochastic, capturing a broader range of behaviors and offering robustness in dynamic environments. Finally, MaxEnt IRL can generalize better to unseen states, making it particularly suitable for complex environments like pedestrian crowds [43]. Lastly, in the low data regime, MaxEnt IRL is comparable, if not superior, to many variants that build on top of the base algorithm using deep neural networks.
Multi-agent Maximum-Entropy IRL [40]. In MaxEnt IRL [40], the goal is to learn a cost function that applies either to a single agent or is designed to collectively model a group of individuals. However, individuals in groups behave differently according to their nature. For instance, some individuals might be more risk averse than others, and therefore, are more likely to have a higher reward for collision avoidance. Others might place greater importance on reaching their goals faster and therefore place greater weight for the cost of distance to goal. Therefore, we include the multi-agent MaxEnt IRL framework proposed by Mehr et al. [40].
We assume for simplicity that a pedestrian’s cost is a weighted linear function $\Phi^{i}\left(\Gamma^{i}_{\theta^{i}}\right)=\sum_{t=0}^{T-1}{\theta^{i}}^{\top}\mathcal{J}^{i}\left(x_{t},\mathbf{u}_{t}\right)$ where $\Phi^{i}\left(\Gamma^{i}_{\theta^{i}}\right)$ computes the features for agent $i$ with respect to weights $\theta^{i}$ and costs $\mathcal{J}^{i}\left(x_{t},\mathbf{u}_{t}\right)$. In this case, solving Equation (1) reduces to finding $\theta^{i,*}$ for every $i$ such that,
@@
Equation (3) can be efficiently solved via block coordinate descent [44] where we treat $\theta$ as a concatenation of each $\theta^{i}$, where $i$ then refers to a coordinate. An update to the current cost parameters, $\theta^{i}_{\textrm{old}}$, of agent $i$ is of the following form,
$\begin{split}\theta^{i}_{\textrm{new}}&\leftarrow\theta^{i}_{\textrm{old}}-\beta\left(\mathbb{E}_{\Gamma^{i}_{\theta^{i}_{\textrm{old}}}\sim\mathcal{D}_{\theta^{i}_{\textrm{old}}}}{\theta^{i}_{\textrm{old}}}^{\top}\Phi^{i}\left(\Gamma_{\theta^{i}_{\textrm{old}}}\right)-\mathbb{E}_{\widehat{\Gamma}^{i}\sim\mathcal{\widehat{D}}}\Phi^{i}\left(\widehat{\Gamma}^{i}\right)\right)\\ \theta^{i}_{\textrm{old}}&\leftarrow\theta^{i}_{\textrm{new}}\\ \end{split}$ (4)
where we compute the features for trajectories generated from the new set of cost parameters $\left(\theta^{i}_{\textrm{new}}\right)$ and proceed to update the cost parameters for the next agent. We iterate over each agent $i$ and repeat this process until convergence. At each iteration of the update process, each agent $i$ uses its parameterized stochastic policy $\mathcal{F}^{i}_{\theta^{i}}\left(x_{t}\right)$ to roll out $M$ trajectories $\{\Gamma^{i}\}_{M}$ corresponding to the updated cost parameters $\left(\theta^{i}_{\textrm{new}}\right)$. In the case of multi-agent IRL, the policy $\mathcal{F}^{i}_{\theta^{i}}$ must be obtained at each iteration using the current parameters.
In multi-agent IRL, the set of policies, $\mathcal{F}^{i}_{\theta^{i}}\left(x_{t}\right)$, for $k$ boundedly rational agents constitutes a (noisy) mixed-Nash solution obtained via backwards recursion through a set of Ricatti equations [37] corresponding to the dynamic Gaussian game $G$ if $G$ is a linear-quadratic game with known linear dynamics and quadratic cost function. In unstructured crowds, however, the pedestrian dynamics are unknown. Recently, Scholler et al. [45] found that constant velocity motion models achieve state of the art results in motion planning tasks on pedestrian trajectory prediction benchmarks. Inspired by this result, we approximate the unknown systems dynamics with constant velocity dynamics,
@@
Furthermore, as we do not know the cost functions $\mathcal{J}^{i}$, we approximate them via Taylor series expansion around the cumulative joint state value $\bar{x}_{t}=x_{1:t}$ to approximate $\mathcal{J}^{i}$ as
$\mathcal{\widetilde{J}}^{i}\left(\bar{x}+\delta_{x_{t}}\right)\approx\mathcal{\widetilde{J}}^{i}\left(\bar{x}\right)+\frac{1}{2}\delta_{x_{t}}^{\top}H_{i}^{t}\delta_{x_{t}}+{l_{i}^{t}}^{\top}\delta_{x_{t}}$ (15)
where $\delta_{x_{t}}=\bar{x}_{t}-x_{t}$ is the difference between the cumulative joint state and the joint state at the current time step. The matrix $H^{i}_{t}$ and vector $l^{i}_{t}$ correspond to the Hessian and gradient of the approximate cost function with respect to $x_{t}$. However, the constant velocity dynamics, combined with the fact that humans typically walk along linear paths at near identical speeds on average, results in a linear approximation of the cost function yielding a non-positive semidefinite $H^{i}_{t}$. This violates the assumptions made in Theorem $2$ of [40] rendering the computation of $\mathcal{F}^{i}_{t}$ intractable. More formally, Theorem $2$ states that $\mathcal{F}^{i}_{t}$ is a Gaussian distribution where the covariance matrix is obtained by computing $\left(H^{i}_{t}\right)^{-1}$, which is not defined for non-PSD matrices.
A solution to achieving tractability consists of modifying the assumption of bounded rationality. More formally, adding a constant factor to the diagonal elements of the learned covariance matrix makes inversion tractable at the cost of approximating the mixed-Nash equilibrium. As seen in our evaluation, this leads to a drop in accuracy in capturing the various behavior modes of individuals walking in a crowd.
IV-B Behavior Cloning [46, 41]
Given a set of expert demonstrations, behavior cloning (BC) aims to directly learn a policy in supervised learning fashion that maps states to actions. For pedestrian crowds, this involves learning how individuals navigate and interact in various scenarios. Expert demonstrations are collected as sequences of state-action pairs over a finite horizon $T$. Using the collected data, we train a model $F^{i}_{\theta^{i}}$ for each agent $i$ to predict the expert’s action given a state. The objective is to minimize the difference between the predicted action and the expert’s action. Formally, the loss $L$ for agent $i$ is given by:
$L(\theta^{i})=\sum_{t=0}^{T}\left\lVert u^{i}_{t}-F^{i}_{\theta^{i}}\left(x^{i}_{t}\right)\right\rVert^{2}$
[…]
$u^{i}_{t}=F^{i}_{\theta^{i}}\left(x^{i}_{t}\right)$
While BC offers a direct approach to imitation learning, it’s essential to acknowledge its inherent limitations in the context of pedestrian crowds. First, BC is purely data-driven and the agent’s behavior is limited to the scope of the expert demonstrations. Small deviations, $\delta x^{i}_{t}=x^{i}_{t}-F^{i}_{\theta^{i}}\left(x^{i}_{t-1}\right)$, from the expert trajectory can therefore lead the agent into states not covered during training, resulting in unpredictable behavior, where $\delta x^{i}_{t}$ represents the deviation at time $t$. This can be problematic in dynamic environments where adaptability is crucial. Second, fewer expert trajectories could result in sub-optimal policies.
The choice of the representative algorithm for Behavior Cloning (BC) depends on the specific neural network architecture employed. For instance, the conventional baseline is the feed-forward neural network, often referred to as the multi-layer perceptron (MLP). Nevertheless, when it comes to capturing the temporal intricacies of trajectories, recurrent neural architectures like LSTMs prove to be a compelling option. Furthermore, the emergence of transformers has marked a significant impact on various computer vision and natural language processing benchmarks, making them another noteworthy candidate for a representative BC algorithm. In this study, we include all of these options as representative algorithms.
Hybrid IRL-BC. BC enables the imitation of expert demonstrations through a supervised learning paradigm. Despite its simplicity and practical results in allowing robots to generalize complex behaviors to new and unstructured scenarios [47], BC often suffers from compounding errors due to the fundamental design of the policy itself. In light of these challenges, we implement a novel hybrid model that synergizes the principles of Inverse Reinforcement Learning (IRL) and BC. This model is designed to uncover the underlying structures and “rewards” guiding expert behavior while preserving the practical and applicable nature of BC. In this innovative approach, we reformulate BC using implicit models, specifically, by employing the composition of $\arg\min$ with a continuous energy function $E_{\theta}$ to represent the policy $\pi_{\theta}$ as:
$u^{i}_{t}=\arg\min_{\mathcal{U}^{i}}E_{\theta}(x_{t},u^{i}_{t})$
This reformulation allows the representation of policies implicitly and formulates imitation as a conditional Energy-Based Modeling (EBM) problem, optimizing actions by minimizing a continuous energy function conditioned on the observations. By integrating the inferred underlying structures of the expert’s behavior, akin to IRL, with the simplicity and direct applicability of BC, this model is considered as a hybrid IRL-BC approach.
IV-C Generative Models [42]
Generative models hold paramount importance in understanding and simulating potential behaviors in agents across multiple domains including computer vision and natural language processing. The overarching objective of these models is to learn and represent behaviors by discerning the underlying distributions and relationships between different states and actions. This is achieved by accurately estimating the parameters—means ($\mu$) and covariances ($\Sigma$)—of fixed Gaussian components. Each Gaussian component in the model represents a potential behavior for the agent, symbolizing unique sets of actions associated with specific states, and is characterized by the probability density function:
$P(x)=\frac{1}{\sqrt{(2\pi)^{k}|\Sigma|}}\exp\left(-\frac{1}{2}(x_{t}-\mu)^{T}\Sigma^{-1}(x_{t}-\mu)\right)$
By learning the statistical properties embedded in these components, generative models are capable of producing actions and data points that are coherent with observed behaviors, thereby allowing for a comprehensive understanding and representation of the variability and diversity inherent in agent actions. Once the model has adequately learned the means and covariances, it predicts actions by sampling from the established distributions using $u^{i}_{t}\sim\mathcal{N}(\mu,\Sigma)$, where $u^{i}_{t}$ is the sampled action for $i$ at time $t$ representing a potential behavior of the agent, influenced by the learned parameters of the distribution. This meticulous process of learning and sampling provides profound insights into the plausible behaviors of agents, enabling the generation of actions that are not only consistent with observed data but also depict the intricate dependencies and variations among different states and actions. In essence, generative models serve as sophisticated tools for capturing and simulating the myriad of possible behaviors in agents by leveraging the relationships and distributions inherent in their actions and states.
(a) Speedway
(b) Waymo/INTERACTION
Figure 3: Visualizing and comparing the entropy of trajectories in the Speedway dataset ($1.061$ bits) with the entropy of trajectories in the Waymo and INTERACTION datasets ($0.336$ and $0.475$ bits, respectively). We observe that trajectories in the Speedway dataset are denser, more unstructured, and have a higher entropy, making it more suited for our analysis.
V Evaluation and Discussion
We compare different IRL, generative, and BC approaches and aim to understand the following questions: $i)$ How do different IL algorithms perform on real world SMGs and $(ii)$ What are the current assumptions needed for successful IL in real world SMGs?
V-A Speedway Dataset
Dataset Has SMGs?
Structured Unstructured Agents involved Density (# agents / SMG)
@@ V-A Speedway Dataset
Waymo [49] ✓(0.336 bits) Vehicles $2$
INTERACTION [50] ✓(0.475 bits) Vehicles $2$
Speedway ✓(1.061 bits) Pedestrians and Vehicles $\mathbf{2-4}$
TABLE I: Comparing of the Speedway dataset with state-of-the-art motion datasets: The Speedway dataset features trajectories of pedestrians as well as vehicles characterized by higher entropy (greater unpredictability) and increased density.
(a) (MA) Max-Ent IRL [40]
@@
We collected $3$ hours of trajectory data at one of the busiest intersections ($24^{\textrm{th}}$ St. and Speedway) on The University of Texas at Austin campus using a Velodyne lidar mounted at the intersection. Our dataset consists primarily of tightly coupled agent interactions that are rarely found in existing autonomous driving and pedestrian datasets such as SCAND [6]. This type of data is useful for building safe and efficient decision-making and motion planning systems for robots in densely populated human environments such as homes, last mile delivery, and autonomous driving.
@@
The lidar continuously streams data as frames where each frame consists of the timestamp and a sequence of frame objects corresponding to agents observed by the lidar. Each frame object consists of a unique id that is persistent in all the frames in which the agent appears, location of the agent in meter with respect to its distance to the sensor, width and length of the bounding box detected for the agent, the angle of the detected bounding box in radian, the class type of the agent such as ‘pedestrian’ or ‘car’, the speed of the agents, and the accuracy of the detection. We recorded trajectories consisting of the id of the agents, their locations, speed, and orientation during busy moments (such as in between classes).

### v2

@@ III-A Background: MaxEnt IRL
The objective in MaxEnt IRL is to adjust the weights of the cost function such that the expected feature counts under the derived policy match the feature counts of the expert, while also maximizing the entropy of the policy. The cost function is iteratively adjusted, typically using gradient-based methods, to ensure that the expected feature counts from the policy induced by the current reward function align with those from the expert demonstrations. Note that since the MaxEnt IRL algorithm traditionally applies to a single agent, Problem (1) is modified to finding $\theta^{*}$ that minimizes
$\left\lVert\mathbb{E}_{\Gamma_{\theta}\sim\mathcal{D}_{\theta}}\Phi\left(\Gamma_{\theta}\right)-\mathbb{E}_{\widehat{\Gamma}\sim\mathcal{\widehat{D}}}\Phi\left(\widehat{\Gamma}\right)\right\rVert$ (3)
There are several reasons why the MaxEnt IRL has been successful. First, the maximum entropy principle ensures a unique solution to the reward recovery problem, mitigating ambiguities inherent in traditional IRL. Further, resultant policies are inherently stochastic, capturing a broader range of behaviors and offering robustness in dynamic environments. Finally, MaxEnt IRL can generalize better to unseen states, making it particularly suitable for complex environments like pedestrian crowds [26].
IV Multi-agent MaxEnt IRL for Unstructued Pedestrian Crowds
In this section, we discuss the technical details of our proposed multi-agent MaxEnt IRL approach. We visually summarize the algorithm in Figure 2. In MaxEnt IRL [26], the goal is to learn a cost function that applies either to a single agent or is designed to collectively model a group of individuals. However, individuals in groups behave differently according to their nature. For instance, some individuals might be more risk averse than others, and therefore, are more likely to have a higher reward for collision avoidance. Others might place greater importance on reaching their goals faster and place greater weight for the cost of distance to goal. Therefore, we propose the following multi-agent formulation of MaxEnt IRL for unstructured pedestrian crowds.
Given a feature function $\Phi^{i}\left(\Gamma^{i}_{\theta^{i}}\right)$,  that computes the features for agent $i$ with respect to weights $\theta^{i}$ and costs $\mathcal{J}^{i}\left(x_{t},\mathbf{u}_{t}\right)$, solving Equation (1) reduces to finding $\theta^{i,*}$ for every $i$ such that,
@@ IV Multi-agent MaxEnt IRL for Unstructued Pedestrian Crowds
Equation (4) can be efficiently solved via block coordinate descent [67] where we treat $\theta$ as a concatenation of each $\theta^{i}$, where $i$ then refers to a coordinate. An update to the cost parameters at the current iteration, $\theta^{i}$, of agent $i$ is of the following form,
$\theta^{i}_{\textrm{new}}\leftarrow\theta^{i}-\beta\left(\mathbb{E}_{\widehat{\Gamma}^{i},\Gamma^{i}_{\theta^{i}}\sim\mathcal{\widehat{D}},\mathcal{D}_{\theta^{i}}}\Phi^{i}\left(\Gamma^{i}_{\theta^{i}}\right)-\Phi^{i}\left(\widehat{\Gamma}^{i}\right)\right)$ (5)
where we compute the features for trajectories generated from the new set of cost parameters $\left(\theta^{i}_{\textrm{new}}\right)$ and proceed to update the cost parameters for the next agent. We iterate over each agent $i$ and repeat this process until convergence. At each iteration of the update process, each agent $i$ uses a dynamic game solver [68] as its parameterized stochastic policy $\mathcal{F}^{i}_{\theta^{i}}\left(x_{t}\right)$ to roll out $M$ trajectories $\{\Gamma^{i}\}_{1:M}$ corresponding to the updated cost parameters $\left(\theta^{i}_{\textrm{new}}\right)$. In the case of multi-agent IRL, the policy $\mathcal{F}^{i}_{\theta^{i}}$ must be obtained at each iteration using the current parameters.
Recently, Mehr et al. [29] showed that when the cost function for an agent, $\mathcal{J}^{i}$, assumes a quadratic form with linear system dynamics, then the set of policies, $\mathcal{F}^{i}_{\theta^{i}}\left(x_{t}\right)$, for $k$ boundedly rational agents can be obtained via backwards recursion through a set of Ricatti equations [36]. In unstructured crowds, however, the pedestrian dynamics are unknown. Recent IRL-based studies found that constant velocity [69] and constant acceleration [26] motion models achieve a good enough approximation for pedestrian trajectory prediction benchmarks. Inspired by this result, we approximate the unknown systems dynamics with constant velocity dynamics,
@@
where $x^{i}_{t}=[p^{i,x},p^{i,y},\psi^{i}]^{\top}$ represents the 2D position and heading angle of agent $i$. Furthermore, as we do not know the cost functions $\mathcal{J}^{i}$, we generate a quadratic approximation via Taylor series expansion around the cumulative joint state value $\bar{x}_{t}=x_{1:t}$ to approximate $\mathcal{J}^{i}$ as
$\mathcal{\widetilde{J}}^{i}\left(\bar{x}+\delta_{x_{t}}\right)\approx\mathcal{\widetilde{J}}^{i}\left(\bar{x}\right)+\frac{1}{2}\delta_{x_{t}}^{\top}H^{i}_{t}\delta_{x_{t}}+{l^{i}_{t}}^{\top}\delta_{x_{t}}$ (16)
where $\delta_{x_{t}}=\bar{x}_{t}-x_{t}$ is the difference between the cumulative joint state and the joint state at the current time step. Note that this formulation only considers nonlinear costs on state variables and the dependence on agents’ actions is quadratic. A similar approximation may be derived for a more general case where the cost function is nonlinear in control as well. The matrix $H^{i}_{t}$ and vector $l^{i}_{t}$ correspond to the Hessian and gradient of the approximate cost function with respect to $x_{t}$. Each policy, $\mathcal{F}^{i}_{\theta^{i}}\left(x_{t}\right)$, is a Gaussian distribution where the mean and covariance parameters are functions of $H^{i}_{t}$:
$\begin{split}\mathcal{F}^{i}_{\theta^{i}}\left(x_{t}\right)&=\mathcal{N}\left(\mu^{i}_{t},\Sigma^{i}_{t}\right)\\ \mu^{i}_{t},\Sigma^{i}_{t}&\leftarrow f\left(H^{i}_{t}\right)^{-1}\\ \end{split}$ (17)
where $f(\cdot)$ is a linear matrix transformation function (c.f. [29] for more details of $f(\cdot)$).
Intractability of Equation (16)
However, the constant velocity dynamics assumption, combined with the fact that humans typically walk along linear paths at near identical speeds on average, results in a linear approximation of the cost function in Equation (16). Under constant velocity dynamics and linear motion, the state evolution is linear. The deviations $\delta_{x_{t}}$ from the trajectory $\bar{x}_{t}$ remain small and linear. Consequently, the second-order derivatives (i.e., the entries of $H_{t}^{i}$) become very small since there is minimal curvature in the cost function along linear trajectories. Mathematically, the curvature of the cost function is captured by the eigenvalues of the Hessian $H_{t}^{i}$. If the cost function has little curvature (as it does under linear motion), these eigenvalues are close to zero, resulting in a non positive semi-definite (PSD) matrix.
Tractability-Rationality Trade-off Trick
To achieve computational tractability, we relax the assumption of bounded rationality, which is typically represented by the stochastic nature of the agent’s policy $\mathcal{F}^{i}_{\theta^{i}}(x_{t})$, characterized by the covariance matrix $\Sigma^{i}_{t}$. The bounded rationality assumption implies that agents exhibit some randomness in their decision-making, often modeled as a probabilistic distribution over actions.
However, performing computations directly with $\Sigma^{i}_{t}$ can lead to intractable solutions due to its dependence on $H_{t}^{i}$ which may be non-PSD. To address this, we modify the covariance matrix by adding a function of its diagonal elements:
$\widetilde{\Sigma}^{i}_{t}=\Sigma^{i}_{t}+g\left(\textrm{diag}(\Sigma^{i}_{t})\right)I,$ (18)
where $g\left(\textrm{diag}(\Sigma^{i}_{t})\right)$ produces a vector that adjusts the diagonal elements of $\Sigma^{i}_{t}$. This operation ensures that $\widetilde{\Sigma}^{i}_{t}$ is a positive semi-definite (PSD) matrix, thus preserving the stochastic nature of the policy while making the problem more tractable. The diagonal elements of $\Sigma^{i}_{t}$, denoted as $\sigma^{ij},i=j$, quantify the uncertainty or randomness in the action that agent $i$ takes toward agent $j$. By applying the function $g$ to these diagonal elements, we effectively increase the variance (uncertainty) of the agent’s actions. This modification shifts towards a more tractable model at the cost of potentially capturing a less accurate representation of agent behaviors, as seen in our evaluation.
Remark: There are several new open problems in this research direction. First, in this work, we heuristically select $g(\cdot)$ from multiple viable candidates $\left(g\left(\textrm{diag}\left(\Sigma^{i}_{t}\right)\right)I=I\right)$ is a trivial choice which enforces tractability but eliminates rationality). Establishing a suitable class of functions for $g(\cdot)$ remains a topic for future work. Another interesting open problem is to establish a bound on the change in the agents’ behavior due to the adjustment of the covariance matrix, $\widetilde{\Sigma}^{i}_{t}$, and its subsequent impact on the mixed-Nash equilibrium.
V Evaluation
We compare our multi-agent MaxEnt IRL approach with MaxEnt IRL, behavior cloning, and a generative modeling approach and aim to understand how our method compares to these baselines in real world pedestrian crowds.
V-A Data Pre-processing
Dataset Description
Structured Unstructured Agents involved Density (# agents / sq. meter)
ETH/UCY [70] - -
@@ V-A Data Pre-processing
Waymo [25] ✓(0.336 bits) Vehicles $1$
INTERACTION [72] ✓(0.475 bits) Vehicles $1$
Speedway ✓(1.061 bits) Pedestrians and Vehicles $\mathbf{3-4}$
TABLE I: Comparing our dataset with state-of-the-art motion datasets: Our dataset features trajectories of pedestrians as well as vehicles characterized by higher entropy (greater unpredictability) and increased density.
Figure 3: Cumulative RMSE distribution demonstrates the percentage of trajectories below a certain RMSE threshold; steeper curves indicate a more effective baseline.
(a) Multi-agent Max-Ent IRL
@@
We collected an unstructured real world pedestrian crowd dataset at one of the busiest intersections on The University of Texas at Austin campus using a BlueCity [^2] https://ouster.com/products/software/blue-city Velodyne LiDAR mounted at the intersection. Our dataset consists of more tightly coupled agent interactions compared to existing motion planning datasets [6, 71, 25]. Specifically, in Figure LABEL:fig:_lines, we visualize the difference between our dataset and an intersection chosen at random from the Waymo dataset, which is considered one of state-of-the-art benchmarks for motion prediction [25]. We observe that trajectories in our dataset are denser and more unstructured (higher entropy), making it an ideal benchmark to test the ability of algorithms to handle unstructured crowds.
@@
The BlueCity LiDAR is equipped with in-built object detection and tracking modules [73]. The LiDAR continuously streams data as frames where each frame consists of the timestamp and a sequence of frame objects corresponding to agents observed by the LiDAR. Each frame object consists of a unique id that is persistent in all the frames in which the agent appears, location of the agent in meter with respect to its distance to the sensor, width and length of the bounding box detected for the agent, the angle of the detected bounding box in radian, the class type of the agent such as ‘pedestrian’ or ‘car’, the speed of the agents, and the accuracy of the detection. We recorded trajectories consisting of the id of the agents, their locations, speed, and orientation during busy moments (such as in between classes).
We processed the data by removing agents that were either at a standstill or appeared at the edges of the LiDAR’s field of view. Furthermore, we clipped each trajectory within a uniform spatio-temporal range ($[-20m,+20m]$ in the x axis and $[-10m,15m]$ in the y axis). After processing the raw trajectories, we ended up with $20$ trajectories of pedestrians with $5$ trajectories each going from east to west (denoted as ‘W’), west to east (denoted as ‘E’), north to south (denoted as ‘S’), south to north (denoted as ‘N’). We created $4$ different categories, ‘W-E-S’,‘W-E-N’,‘S-N-W’,‘S-N-E’, with $1255^{3}=$ 125 trajectories each for a total of $500$ trajectories.